OpenStudy (anonymous):

Solve for the variable. 1/5p=14

OpenStudy (anonymous):

5p*14=70p...(1/70)=p

OpenStudy (anonymous):

unless you meant (1/5)*p = 14....

OpenStudy (anonymous):

then p=70.

OpenStudy (anonymous):

a) (x + 11)(x + 4) b) (x + 11)(x - 4) c) (x - 11)(x - 4) d) (x - 11)(x + 4)

OpenStudy (anonymous):

my bad